### v2

@@ Abstract
Ship trajectories from Automatic Identification System (AIS) messages are important in maritime safety, domain awareness, and algorithmic testing. Although the specifications for transmitting and receiving AIS messages are fixed, it is well known that technical inaccuracies and lacking seafarer compliance lead to severe data quality impairment. This paper proposes an adaptable, data-driven, $\alpha$-quantile-based framework for decoding, constructing, splitting, and assessing trajectories from raw AIS records to improve transparency in AIS data mining. Results indicate the proposed filtering algorithm robustly extracts clean, long, and uninterrupted trajectories for further processing. An open-source Python implementation of the framework is provided.
@@ 1 Introduction and background
Concerning the methodology for trajectory extraction, a notable observation emerges. While a confluence of approaches exists, an agreed-upon standard within the scientific community is still absent. Trajectory extraction usually consists of two consecutive steps. First, filtering the data set on an individual message level to eliminate erroneous messages. Second, the cleaned messages are combined into meaningful, coherent trajectories.
@@
Following the initial filtration of individual messages, trajectories are typically constructed by sequentially linking messages emanating from the same source, identified by the Maritime Mobile Service Identity (MMSI) (Mao et al., 2018; Zhao et al., 2018; Yuan et al., 2019; Guo et al., 2021). Subsequently, an additional layer of trajectory-based filtering is applied to the generated trajectories. Several criteria have been proposed for filtering and splitting trajectories to enhance the accuracy and granularity of maritime data. Yuan et al. (2019), Zhao et al. (2018), and Mao et al. (2018) have contributed specific conditions for these purposes, while Yuan et al. (2019) suggests filtering trajectories based on the number of observations, proposing a threshold of 20 observations per track. Zhao et al. (2018) introduces a more stringent criterion, recommending a minimum of 100 observations for a track to be considered, whereas Mao et al. (2018) sets a threshold of 500 observations for tracks where the $SOG$ is not equal to zero, thereby emphasizing the importance of trajectories during active movement. Additionally, Mao et al. (2018) proposes a condition to avoid high-complexity trajectories, employing the average cosine of the angles between three consecutive message points as an indicator, with a threshold set at $0.8$.
Moreover, the concept of splitting trajectories into sub-tracks is addressed by Zhao et al. (2018). This involves three key conditions: a speed change greater than 15 knots between consecutive messages, a time interval exceeding 10 minutes between messages, and a permissible deviation between reported and generated time within $[-5,5]$ seconds, covering 96% of the ranges. These conditions collectively provide a comprehensive approach to delineating sub-tracks within longer trajectories, thereby extracting meaningful and discrete segments of vessel movement from AIS records.
A shared challenge in the aforementioned revised studies lies in determining diverse thresholds crucial for filtering records or trajectories. While certain thresholds are derived from empirical data, the majority are expert-derived, raising concerns about their generalizability and justification. In this paper, we advocate for an alternative approach to trajectory extraction, characterized by its reliance on data-driven methodologies with minimal reliance on predefined threshold values. Recognizing the inherent limitations and subjectivity associated with conventional threshold-setting, we aim to redefine trajectory extraction by minimizing assumptions and embracing a more empirically driven paradigm. Central to our approach is utilizing the $\alpha$-quantiles of empirical distributions constructed from different variables in the data, whose values will be used to distinguish coherent from incoherent trajectories.
Consequently, a comprehensive analysis was conducted on a dataset spanning 912 days of AIS records from January 2020 to June 2022. The geographical scope of this investigation encompasses the North Sea and regions of the Baltic Sea, with the primary objective of determining robust methodologies for extracting long, uninterrupted, and clean trajectories from the dataset. By clean, we understand the absence of erratic and highly volatile movement sequences of navigational parameters (SOG, COG, Position). All methodology presented in this article is publicly available as an open-source package (Paulig, 2024) for the Python programming language.
The rest of the paper is organized as follows. Section 2 briefly revisits AIS records and their structure, and Section 3 details the obtained data set. Section 4 introduces the split-point procedure based on empirical distribution functions, whose properties are detailed in Section 5. Section 6 proposes tools for assessing spatial features of the extracted trajectories, while Section 7 discusses the obtained results. Section 8 concludes. Additionally, C provides exemplary usage of the accompanying Python package.
@@ 3 AIS data set
Figure 1: Geographical extent of the data set
@@
The decoded AIS messages, however, cannot be used directly for trajectory extraction as they contain notable inconsistencies (Bailey, 2005; Harati-Mokhtari et al., 2007). Therefore, individual messages are analyzed to identify and eliminate those with erroneous positional and speed data. Subsequently, these messages are aggregated to construct trajectories. Given that a single trajectory may comprise multiple sub-trajectories, a data-driven, rule-based system is used to segment them.
@@ 3.2 Velocity reports
Figure 4: Trajectory of the vessel with MMSI 211XXXX90 laying at anchor in the harbor of Emden, Germany.
@@ 4.2 Trajectory splitting
Consistent with the abovementioned research, we use the MMSI to assign AIS messages to their originating vessels and order these messages using the UTC timestamp to formulate a trajectory. Nevertheless, this methodology exhibits significant shortcomings without further processing, as it often results in trajectories with numerous undesirable characteristics, which will be outlined in this section. Accordingly, this study advances the trajectory segmentation methodology proposed by Zhao et al. (2018), applying it to ascertain the continuity of a trajectory by evaluating whether two successive AIS messages should be attributed to the same navigational path via a split-point approach. Therefore, $\alpha$-quantiles of various distributions, detailed in the next Sections, are computed to decide whether to split the trajectory at a particular point. The procedure is consolidated in Algorithm 1. Throughout the next subsections, it is assumed that if a trajectory is split multiple times such that any of the sub-trajectories only consists of one message, this sub-trajectory is discarded.
@@ 4.2.1 Temporal difference
The most natural metric to investigate for split points is the temporal difference between two successive messages of one trajectory $\operatorname{unix}(m_{i+1})-\operatorname{unix}(m_{i})$. Revisiting the AIS specifications, the defined transmission interval for dynamic AIS messages ranges from two to ten seconds for class A equipment ($2-30$ seconds for class B equipment) for vessels not at anchor and 180 seconds for vessels at anchor. Therefore, it is expected to find an accumulation of temporal gaps around these thresholds in the data, which could then be used to split trajectories. To verify the assumptions, a histogram of the temporal differences between two successive messages inside each trajectory for July 2021 is depicted in Figure 5. A brief discussion on why only a single month of data was used in the calculation is provided in Section 5. The distribution clearly shows that our initial assumptions are untenable, as temporal differences vary significantly in the interval from $[2,360]s$ with an exceptional peak around the $360s$ mark, whose origin remains undisclosed. The only message types in the AIS specification transmitted at $360s$ intervals are safety-relevant messages, none in the data set, and static voyage reports, which had not been considered.
The findings of this study suggest that it is not advisable to segment vessel trajectories based on the transmission intervals prescribed by AIS standards. Exemplary, we propose using the $(1-\alpha)\cdot 100=95$ percentile of the empirical distribution of temporal gaps between AIS messages to identify temporal split points. Following this logic, a split-point in this data set is given when the temporal interval between two successive messages exceeds $392s$ (see also Figure 5).
@@ 4.2.2 Velocity changes
Figure 6: Histograms absolute SOG first differences (left), and turning rates for July 2021.
Following the procedure from the preceding paragraph, the absolute speed change distribution across all vessel trajectories for July 2021 was obtained. The results are illustrated on the right side of Figure 6, showing that only $\alpha=5\%$ of speed changes in individual trajectories are larger than $2.6kn$, resulting in a possible acceleration/deceleration range from $\frac{2.6\,kn}{392s}=0.007\,kn/s$ to $\frac{2.6\,kn}{2s}=1.3\,kn/s$, which are reasonable values, even for larger ships. This study identifies two messages with more than $2.6kn$ of absolute speed change as a split point, cutting it into two sub-tracks at this position.
@@ 4.2.3 Turning rates
It is generally not anticipated that significant deviations in a ship’s turning rate will occur during navigation, as rapid directional shifts not only increase material wear and tear but also contravene the IMO’s routing rationale advocating for safe and straightforward routing (IMO, 2019). Consequently, the right side of Figure 6 presents a histogram of tuning rates for two consecutive messages from one trajectory for July 2021, which will be used for threshold determination. Turning rates are calculated by dividing the difference of COG changes of two consecutive messages by their temporal difference. Adhering to the same $(1-\alpha)\cdot 100=95$ percentile criterion as applied to SOG changes, the eligible range of turning rates between two messages lies in the interval $[-0.48^{\circ}/s,0.38^{\circ}/s]$.
@@ 4.2.4 Reported against positional velocity
Another deficiency of trajectories that cannot be captured with the three metrics above are erroneous position reports that lead to significant outliers, as seen in Figure 8. In those cases, there is usually no anomaly detectable in terms of speed, course, or temporal differences, which is why this study follows Zhao et al. (2018) and uses the difference between $\mkern 1.5mu\overline{\mkern-1.5muSOG\mkern-1.5mu}\mkern 1.5mu_{m_{i}}^{m_{i+1}}$ and $\widehat{SOG}_{m_{i}}^{m_{i+1}}$ to determine the outliers. Figure 9(a) provides context on the magnitude of differences across the admitted speed range. It can be observed that there are substantial deviations of different magnitudes in both directions. As with the other metrics, an empirical distribution of the differences is depicted in Figure 9(b) to establish a data-driven threshold. The $(1-\alpha)\cdot 100=95$ percentile for the present data falls within the left-skewed interval $[-8.96,6.65]\,kn$. The skewness of the distribution is expected, as $\mkern 1.5mu\overline{\mkern-1.5muSOG\mkern-1.5mu}\mkern 1.5mu_{m_{i}}^{m_{i+1}}$ is bounded at $30kn$ due to the speed filtering employed in Section 3.2, while $\widehat{SOG}_{m_{i}}^{m_{i+1}}$ is unbounded.
(a) Scatter plot of the average observed SOG ($\mkern 1.5mu\overline{\mkern-1.5muSOG\mkern-1.5mu}\mkern 1.5mu_{m_{i}}^{m_{i+1}}$) and the calculated SOG from positional and temporal data ($\widehat{SOG}_{m_{i}}^{m_{i+1}})$ for all trajectories of the month of July 2021
(b) Histogram of the distribution of $\mkern 1.5mu\overline{\mkern-1.5muSOG\mkern-1.5mu}\mkern 1.5mu_{m_{i}}^{m_{i+1}}-\widehat{SOG}_{m_{i}}^{m_{i+1}}$
Figure 9: Details on the divergence of $\mkern 1.5mu\overline{\mkern-1.5muSOG\mkern-1.5mu}\mkern 1.5mu_{m_{i}}^{m_{i+1}}$ and $\widehat{SOG}_{m_{i}}^{m_{i+1}}$

### v3

@@ Abstract
Ship trajectories from Automatic Identification System (AIS) messages are important in maritime safety, domain awareness, and algorithmic testing. Although the specifications for transmitting and receiving AIS messages are fixed, it is well known that technical inaccuracies and lacking seafarer compliance lead to severe data quality impairment. This paper proposes an adaptable, data-driven, maneuverability-dependent, $\alpha$-quantile-based framework for decoding, constructing, splitting, and assessing trajectories from raw AIS records to improve transparency in AIS data mining. Results indicate the proposed filtering algorithm robustly extracts clean, long, and uninterrupted trajectories for further processing. An open-source Python implementation of the framework is provided.
@@ 1 Introduction and background
Trajectories, however, cannot be constructed directly from raw AIS messages as they contain notable inconsistencies (Bailey, 2005; Harati-Mokhtari et al., 2007) such as erroneous positions, time, speed or course, Maritime Mobile Service Identity (MMSI) misuse or sharing or intentional transceiver deactivation, rendering them useless for further analysis. Trajectory extraction usually consists of two consecutive steps. First, filtering the raw data on an individual message level to eliminate erroneous messages. Second, the cleaned messages are combined into meaningful, coherent trajectories. A shared challenge for these steps lies in determining the diverse thresholds crucial for filtering individual messages or entire trajectories, especially since the range of many of the possible parameters to consider, like velocity, acceleration or turning rate depends on the maneuvering capabilities of individual ships, of which many different share the national and international waters.
In this paper, we introduce a framework (Figure 1) for maneuverability-dependent trajectory extraction, characterized by its near-complete reliance on data-driven methodologies with minimal reliance on predefined threshold values. The framework aims to cover the entire process of trajectory extraction from raw AIS messages, beginning with the collection and decoding phase (marked with green in Figure 1), MMSI grouping and individual message exclusion (marked with Purple), trajectory and sub-trajectory determination (Blue), and post-determination refinement (Yellow). In the following, relevant section headings will be marked with their corresponding framework color. In particular, our work contributes to the field as follows:
  • Generalizability: Due to the pure data-driven approach, the framework is universally usable, regardless of the geographical region under consideration.
  • Expanded methodology: This article goes beyond the methodology of Zhao et al. (2018), which splits trajectories into sub-tracks based on velocity and time differences between individual messages by additionally considering turning rates, distances, and the maneuverability of the ship by using its length as a proxy.
  • Trajectory assessment: We introduce the concept of average absolute change of course, which allows for flexible assessment of extracted trajectories in terms of maneuvering behavior.
  • Implementation: Our end-to-end framework is publicly available as an open-source package (Paulig, 2024) for the Python programming language.
Central to our approach is utilizing the $\alpha$-quantiles of empirical distributions constructed from different features in the data, whose values will be used to distinguish coherent from incoherent trajectories. The parameter $\alpha$ can be adjusted to control the statistical power of the extracted trajectories.
To test the framework, a comprehensive analysis was conducted on a dataset spanning 912 days of AIS records from January 2020 to June 2022. The geographical scope of this investigation encompasses the North Sea and regions of the Baltic Sea, with the primary objective of determining robust methodologies for extracting long, uninterrupted, and clean trajectories from the dataset. By clean, we understand the absence of erratic and highly volatile movement sequences of navigational parameters (SOG, COG, Position). Our experiment shows that the fully data-driven paradigm is suitable to effectively remove erroneous messages and construct meaningful trajectories out of the remaining messages.
The rest of the paper is organized as follows. Section 2 recapitulates the different methodologies in trajectory extraction so far, Section 3 briefly revisits AIS records and their structure, and Section 4 details the obtained data set and Section 5 explains the individual message exclusion processes. Applications and comparisons of the framework are conducted in Sections 7 and 8. Section 6 introduces the split-point procedure based on empirical distribution functions. Section 9 proposes tools for assessing spatial features of the extracted trajectories, while Section 10 discusses the obtained results. Section 11 concludes. Additionally, Appendix C provides exemplary usage of the accompanying Python package.
Figure 1: Flow chart diagram of the trajectory extraction framework presented in this article. Data collection and decoding is discussed in Section 4, Filtering in Section 5, Determination in Section 6 and post-filtering in Section 9.
2 Related works
Concerning the methodology for trajectory extraction, a notable observation emerges. While a confluence of approaches exists, an agreed-upon standard within the scientific community is still absent.
@@ 2 Related works
Following the initial filtration of individual messages, trajectories are typically constructed by sequentially linking messages emanating from the same source, identified by the MMSI (Mao et al., 2018; Zhao et al., 2018; Yuan et al., 2019; Guo et al., 2021; Capobianco et al., 2021). Subsequently, an additional layer of trajectory-based filtering is applied to the generated trajectories. Several criteria have been proposed for filtering and splitting trajectories to enhance the accuracy and granularity of maritime data. Yuan et al. (2019), Zhao et al. (2018), and Mao et al. (2018) have contributed specific conditions for these purposes, while Yuan et al. (2019) suggests filtering trajectories based on the number of observations, proposing a threshold of 20 observations per track. Zhao et al. (2018) introduces a more stringent criterion, recommending a minimum of 100 observations for a track to be considered, whereas Mao et al. (2018) sets a threshold of 500 observations for tracks where the $SOG$ is not equal to zero, thereby emphasizing the importance of trajectories during active movement. Additionally, Mao et al. (2018) proposes a condition to avoid high-complexity trajectories, employing the average cosine of the angles between three consecutive message points as an indicator, with a threshold set at $0.8$.
Moreover, the concept of splitting trajectories into sub-tracks is addressed by Zhao et al. (2018). This involves three key conditions: a speed change greater than 15 knots between consecutive messages, a time interval exceeding 10 minutes between messages, and a permissible deviation between reported and generated time within $[-5,5]$ seconds, covering 96% of the value range. These conditions collectively provide a comprehensive approach to delineating sub-tracks within longer trajectories, thereby extracting meaningful and discrete segments of vessel movement from AIS records.
@@ 4 Raw data and decoding ( )
Figure 2: Geographical extent of the data set.
@@
5 Individual message exclusion ( )
@@ 5.2 Velocity reports
Figure 5: Trajectory of the vessel with MMSI 211XXXX90 laying at anchor in the harbor of Wismar, Germany.
@@ 6.2 Trajectory splitting
Consistent with the abovementioned research, we use the MMSI to assign AIS messages to their originating vessels and order these messages using the UTC timestamp to formulate a trajectory. Nevertheless, this methodology exhibits significant shortcomings without further processing, as it often results in trajectories with numerous undesirable characteristics, which will be outlined in this section. Accordingly, this study advances the trajectory segmentation methodology proposed by Zhao et al. (2018), applying it to ascertain the continuity of a trajectory by evaluating whether two successive AIS messages should be attributed to the same navigational path via a split-point approach. Therefore, $\alpha$-quantiles of various distributions, detailed in the next Sections, are computed to decide whether to split the trajectory at a particular point. Additionally, we incorporate the ideas from Zhang et al. (2018), and Chen et al. (2020) by explicitly taking the vessel maneuverability into account, by using the ship length as a proxy. Preliminary testing concluded that other proxies such as the ship type work worse, as there are ship types such as CARGO spanning the entire range of ship lengths and widths, therefore not representing maneuverability accurately. Ship mass, while being a promising proxy, was also discarded, as critical information for its calculation are either unreliable in the data (draft) or not part of it (block coefficient). Furthermore, Hansen et al. (2022) related ship lentgths to thier heading change per meter and concluded that ship length is a valid proxy for maneuverability. To get a useful segmentation of the length, we introduce 9 distinct bins $[0,25,50,75,100,125,150,175,200,\infty]m$. We acknowledge, that these boundaries have no physical justification, but are a practical tradeoff between accuracy and computational complexity. Empirical distributions are then calculated for each of the bins. The procedure is consolidated in Algorithm 1 and a visual representation of the length-dependent quantile values can be found in Figure 6. Throughout the next subsections, it is assumed that if a trajectory is split multiple times such that any of the sub-trajectories only consists of one message, this sub-trajectory is discarded.
Figure 6: Length-dependent quantile values for the five different metrics used to determine split points in the constructed trajectories. Note that the time difference quantiles are not used ship-length dependent, and are only displayed for completeness.
@@ 6.2.1 Temporal difference
The most natural metric to investigate for split points is the temporal difference between two successive messages of one trajectory $\operatorname{unix}(m_{i+1})-\operatorname{unix}(m_{i})$. Revisiting the AIS specifications, the defined transmission interval for dynamic AIS messages ranges from two to ten seconds for class A equipment ($2-30$ seconds for class B equipment) for vessels not at anchor and 180 seconds for vessels at anchor. Therefore, it is expected to find an accumulation of temporal gaps around these thresholds in the data, which could then be used to split trajectories. A histogram of the temporal differences between two successive messages inside each trajectory for 2021 is depicted in Figure 7 to verify the assumptions. A brief discussion on why one year of data was used in the calculation is provided in Appendix B. It is noteworthy that we abstained from determining the distribution for each length interval, as the temporal distribution is not influenced by the ship’s length but rather by technical or human factors such as delay or intent. The distribution clearly shows that our initial assumptions are untenable, as temporal differences vary significantly in the interval from $[2,360]s$ with an exceptional peak around the $360s$ mark, whose origin remains undisclosed. The only message types in the AIS specification transmitted at $360s$ intervals are safety-relevant messages, none in the data set, and static voyage reports, which had not been considered.
The findings of this study suggest that it is not advisable to segment vessel trajectories based on the transmission intervals prescribed by AIS standards. Instead the calculated values from the distribution of Figure 7 are used, depending on $\alpha$.
To determine all following empirical distribution functions, we only consider the value of the respective metric if the two messages under consideration are less than the 95%-percentile of temporal gaps, namely $391s$, apart. We do this to explicitly exclude artificial outliers that are so far apart temporally that the significance of the obtained values is questionable.
@@ 6.2.2 Velocity changes
Following the procedure from the preceding paragraph, the absolute speed change distribution for all vessel length intervals for 2021 was obtained. The results are illustrated in Table 3 and show a clear pattern of receding speed and thus acceleration values for increasing ship lengths, reflecting the different maneuvering capabilities of different ship types.
Length bin $\alpha$ Accel. range at $\alpha=0.05$ Length bin $\alpha$ Accel. range at $\alpha=0.05$ Length bin $\alpha$ Accel. range at $\alpha=0.05$
0.1 0.05 0.01 0.1 0.05 0.01 0.1 0.05 0.01
$[0,25)m$ 2.3 4.7 14.4 $[0.012,2.350]$ $[75,100)m$ 0.4 0.9 3.4 $[0.002,0.450]$ $[150,175)m$ 0.4 0.8 2.8 $[0.002,0.400]$
$[25,50)m$ 1.9 3.7 9.0 $[0.009,1.850]$ $[100,125)m$ 0.4 0.9 3.5 $[0.002,0.450]$ $[175,200)m$ 0.3 0.5 1.8 $[0.001,0.250]$
$[50,75)m$ 1.2 2.5 6.2 $[0.006,1.250]$ $[125,150)m$ 0.5 1.3 5.9 $[0.003,0.650]$ $[200,\infty)m$ 0.3 0.6 1.9 $[0.002,0.300]$
Table 3: Cutoff quantile values for the speed of two consecutive messages in miles for selected values of $\alpha$ across the different length intervals. The lower acceleration values are calculated by dividing the cutoff value by the $5\%$-percentile of the temporal difference distribution ($391s$). The upper acceleration values are obtained by dividing by the minimum AIS sending frequency of $2s$. Units of acceleration are $kn/s$.
@@ 6.2.3 Turning rates
It is generally not anticipated that significant deviations in a ship’s turning rate will occur during navigation, as rapid directional shifts as those in Figure 8 not only increase material wear and tear but also contravene the IMO’s routing rationale advocating for safe and straightforward routing (IMO, 2019). Consequently, Table 4 presents the tuning rates for two consecutive messages using all trajectories from 2021 and three selected $\alpha$ values. Turning rates are calculated by dividing the difference of COG changes of two consecutive messages by their temporal difference. As with the differences in speed, we also see a clear length-dependent pattern of inversely related turning rates and ship lengths.
@@
Length bin $[1-\alpha/2,\alpha/2]$ Length bin $[1-\alpha/2,\alpha/2]$
0.1 0.05 0.01 0.1 0.05 0.01
$[0,25)m$ $[-0.619,0.473]$ $[-1.067,0.840]$ $[-1.949,1.600]$ $[75,100)m$ $[-0.113,0.106]$ $[-0.218,0.200]$ $[-0.436,0.408]$
$[25,50)m$ $[-0.521,0.402]$ $[-0.880,0.700]$ $[-1.557,1.300]$ $[100,125)m$ $[-0.100,0.092]$ $[-0.183,0.169]$ $[-0.365,0.340]$
$[50,75)m$ $[-0.305,0.251]$ $[-0.510,0.459]$ $[-0.860,0.785]$ $[125,150)m$ $[-0.099,0.091]$ $[-0.173,0.160]$ $[-0.324,0.300]$
Length bin $[1-\alpha/2,\alpha/2]$
0.1 0.05 0.01
$[150,175)m$ $[-0.083,0.080]$ $[-0.148,0.142]$ $[-0.148,0.142]$
$[175,200)m$ $[-0.062,0.060]$ $[-0.114,0.111]$ $[-0.218,0.210]$
$[200,\infty)m$ $[-0.065,0.062]$ $[-0.113,0.109]$ $[-0.205,0.200]$
Table 4: Cutoff quantile intervals for the turning rate of two consecutive messages in $[{}^{\circ}/s]$ for selected values of $\alpha$ across the different length intervals.
@@ 6.2.4 Reported against positional velocity
Another deficiency of trajectories that cannot be captured with the three metrics above are erroneous position reports that lead to significant outliers, as seen in Figure 9. In those cases, there is usually no anomaly detectable in terms of speed, course, or temporal differences, which is why this study follows Zhao et al. (2018) and uses the difference between $\mkern 1.5mu\overline{\mkern-1.5muSOG\mkern-1.5mu}\mkern 1.5mu_{m_{i}}^{m_{i+1}}$ and $\widehat{SOG}_{m_{i}}^{m_{i+1}}$ to determine the outliers. Figure 10 provides context on the magnitude of differences across the admitted speed range. It can be observed that there are substantial deviations of different magnitudes in both directions. As with the other metrics, some of the determined threshold values can be viewed in Table 5. The skewness of the intervals is expected, as $\mkern 1.5mu\overline{\mkern-1.5muSOG\mkern-1.5mu}\mkern 1.5mu_{m_{i}}^{m_{i+1}}$ is bounded at $30kn$ due to the speed filtering employed in Section 5.2, while $\widehat{SOG}_{m_{i}}^{m_{i+1}}$ is unbounded.
Length bin $[1-\alpha/2,\alpha/2]$ Length bin $[1-\alpha/2,\alpha/2]$
0.1 0.05 0.01 0.1 0.05 0.01
$[0,25)m$ $[-1.434,2.332]$ $[-3.237,4.022]$ $[-8.412,6.699]$ $[75,100)m$ $[-1.074,0.822]$ $[-2.701,1.507]$ $[-10.197,3.338]$
$[25,50)m$ $[-1.426,1.806]$ $[-3.233,3.240]$ $[-9.794,5.381]$ $[100,125)m$ $[-1.294,0.975]$ $[-3.353,1.861]$ $[-12.804,4.014]$
$[50,75)m$ $[-1.032,0.995]$ $[-2.110,1.876]$ $[-6.892,3.633]$ $[125,150)m$ $[-1.294,0.987]$ $[-3.041,1.948]$ $[-11.222,4.233]$
Length bin $[1-\alpha/2,\alpha/2]$
0.1 0.05 0.01
$[150,175)m$ $[-2.023,1.393]$ $[-5.134,2.780]$ $[-19.192,5.877]$
$[175,200)m$ $[-1.534,1.188]$ $[-4.032,2.433]$ $[-15.734,5.155]$
$[200,\infty)m$ $[-2.176,1.371]$ $[-5.907,2.850]$ $[-21.707,6.044]$
Table 5: Cutoff quantile intervals for the difference between reported and calculated speed of two consecutive messages in $[kn]$ for selected values of $\alpha$ across the different length intervals.
Figure 10: Scatter plot of the average observed SOG ($\mkern 1.5mu\overline{\mkern-1.5muSOG\mkern-1.5mu}\mkern 1.5mu_{m_{i}}^{m_{i+1}}$) and the calculated SOG from positional and temporal data ($\widehat{SOG}_{m_{i}}^{m_{i+1}})$ for all trajectories of July 2021.